An ellipse having foci at (3,3) &(-4,4) &passing through the origin has eccentricity equql to

Grade:12th pass

2 Answers

Arun
4 years ago
Since origin in on the elipse therefore assum P(0,0) on elipse,Let S1 = (3,3) and S2 = (-4,4)Since we know that, PS1 + PS2 =2aor2a = 7root(2) – --(1)also S1S2 = 2aeor 2ae = 5 root(2) --- (2)Divide (2) by (1) and get e = 5/7
Vikas TU
4 years ago
Dear student 
P(3,3)
Q(-4,4)
O(0,0)
e = PQ/(OP+OQ)
e = sqrt[(3-4)^2 + (3+4)^2] / sqrt(9+9) + sqrt(16+16)
e = sqrt 50 / 7root 2
= 5 root 2 / 7 root 2 
= 5/7